OpenStudy (anonymous):

which of the following sentence is thrue about the graphs of y=3(3-5)^2+1 and y=3(x+5)^2+1?

OpenStudy (anonymous):

Possible answers:Their vertices are maximums. One graph has a vertex that is a maximum, while the other have a vertex that is a minimum. The graphs have different shapes with different vertices. The graphs have the same shape with different vertices.

OpenStudy (mathmale):

By y=3(3-5)^2+1, don't you really mean y=3(x-5)^2+1? (x, not 3)

OpenStudy (mathmale):

@monkeyme25 : I appreciate your trying to help, but OpenStudy rules request that you guide the other person towards finding the solution and also ask that you DO NO PROVIDE ANSWERS. Please stop (unless you can guide without giving answers).

OpenStudy (mathmale):

@nojelis40 " Which of the following sentence is thrue about the graphs of y=3(x-5)^2+1 and y=3(x+5)^2+1? (Note that I've re-written your (3-5) as (x-5).